OpenStudy (anonymous):

What is the sum of a 22–term arithmetic sequence where the first term is 55 and the last term is –29? please helpp

OpenStudy (cwrw238):

sum of n terms = (n/2)[ a + l] where n = number of terms, a = first term and l = last term

OpenStudy (cwrw238):

so plug in your values to get sum of 22 terms = (22/2)[ 55 - 29]

OpenStudy (anonymous):

i got 286, i keep forgetting the different formulas

OpenStudy (cwrw238):

286 is correct